### v1

# Magnetic properties of Cr8 and V8 molecular rings from ab initio calculations

Abstract
Molecular nanomagnets are systems with a vast phenomenology and are very promising for a variety of technological applications, most notably spintronics and quantum information. Their low-energy spectrum and magnetic properties can be modeled using effective spin Hamiltonians, once the exchange coupling parameters between the localized magnetic moments are determined. In this work we employ density functional theory (DFT) to compute the exchange parameters between the atomic spins for two representative ring-shaped molecules containing eight transition-metal magnetic ions: Cr8 and V8. Considering a set of properly chosen spin configurations and mapping their DFT energies on the corresponding expressions from a Heisenberg Hamiltonian, we compute the exchange couplings between magnetic ions which are first, second and further neighbors on the rings. In spite of their chemical and structural similarities the two systems exhibit very different ground states: antiferromagnetic for Cr8, ferromagnetic for V8, which also features non-negligible couplings between second nearest neighbors. A rationalization of these results is proposed that is based on a multi-band Hubbard model with less-than-half filled shells on magnetic ions.
Molecular Nano-Magnets, DFT+$U$, DFT+$U$+$V$, spintronics, quantum information
I Introduction
Figure 1: Cr8/V8 molecular structure. Colors code: Cr/V green, F orange, O red, C grey, H white.
Figure 2: 3D molecular structure highlighting the anion octahedra around each TM ion.
Molecular nanomagnets (MNM) are attracting a vast interest in the scientific community, due to their huge potential for technological applications [1], including spintronics [2, 3, 4, 5, 6], magneto-refrigeration [7, 8, 9, 10], miniaturisation of data-storage media [11], quantum sensing [12] and quantum-information [13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23]. The reduced size of these systems and, in many cases, the availability of experimental data for comparison, make them ideal to even investigate fundamental questions such as the different behavior of integer and half-integer spin systems [24, 25, 26, 27, 28, 29, 30, 31], the influence of geometrical curvature on the spin order and magnetic interactions [32, 33], the origin of generalized spin-orbit couplings [34], the relation between spin and transport properties [35].
Here, we consider a specific class of MNM: ring-shaped molecules with eight magnetic centers (transition-metal - TM - ions with localized spins) occupied by either Cr or V, known as Cr8 and V8, respectively. Due to their modular and stable structure (TM-centered anion octahedra glued together by organic ligands) and the interchangeable TM ions, they offer a platform to control a variety of magnetic behaviors, which could be engineered for target applications. In fact, these systems have been studied for over two decades, both experimentally [36, 37, 38, 39, 40] and theoretically [41, 42, 43], with particular attention to Cr-based rings, due to their potential for quantum information technologies [44, 45, 46, 47, 48, 49, 50].
Cr8 is the prototype of this class of systems and has been considered both in its pristine and doped compositions with one Cr atom substituted by other TMs to achieve a finite total spin or other interesting magnetic properties [51, 43, 52, 53, 54, 55, 56, 57, 58, 59, 60, 38]. Cr8 features an antiferromagnetic (AFM) order across the ring, with the Cr half integer spins ($S=3/2$) aligned perpendicularly to the molecular plane, in a collinear fashion. This order is quite common among eight-center rings based on 3d magnetic centers.
In contrast, octanuclear rings based on V, like V8, have been studied sporadically in literature [61, 62] and never with the specific structure/composition we investigate here, due to its instability. The only studies we are aware of on V-based rings are on doped magnetic wheels of the form V7X, where X is a TM different from V [39]). Thus, to the best of our knowledge, this work is the first to address the behavior of V8. Due to the 3+ oxidation state of the TM ions, V8 features an integer $S=1$ spin on each magnetic center, consistent with the V ion having one less 3d electron than Cr. However, at variance with similar octanuclear rings based on Cr or other 3d ions, we find V8 to exhibit a peculiar ferromagnetic (FM) behaviour, also characterized by non-negligible AFM couplings between second nearest neighbor magnetic ions.
In this work, we present a first computational study of V8 aimed at investigating, in comparison with the well-known Cr8, its peculiar ground state magnetic properties. Density functional theory [63, 64] calculations are used to determine the electronic ground state of both systems and compare the energy of various spin configurations. Precise values of the exchange couplings between the TM ions are obtained from DFT, upon fixing the sign (i.e., the direction) of the spin on each magnetic site, while their magnitude remains unchanged.
The paper is organized as follows: in Section II we introduce the Cr8 and V8 molecular structures. In Section III we briefly outline the theory of spin Hamiltonians and their integration with DFT calculations. The results of our magnetic couplings calculations are presented in Section IV, while in Section V we discuss our main findings, giving a rationale of the various approximations used. Section VI concludes and summarizes the work.
II Cr8 and V8 molecular complexes
[…]
As mentioned earlier, Cr8 is a well-known system, which has been synthesized and studied thoroughly (even in presence of TM substitutional doping) both experimentally [38, 57, 60, 58, 55], and theoretically [43, 59, 52, 53, 54, 56, 56] via DFT with hybrid functionals, Hubbard corrections, or in conjunction with many-body models.
Unlike Cr8, the specific V8 ring we consider in this work has never been synthesized, due to parasitic oxidation processes [39]. Therefore, a direct comparison of our results with experiments is currently not feasible. To the best of our knowledge, the only two homometallic rings containing eight V magnetic centers ever synthesized and characterized experimentally are reported in Refs. [61] and [62]. However, compared to our V8, they feature a different organic scaffold around the TM centers, and, quite interestingly, report different ground-state magnetic configurations (AFM and FM, respectively). On the other hand, two heterometallic V-based rings with the same structure as our V8 (see Fig. 2) have been synthesized, with single Ni and Zn impurities (V7Ni and V7Zn, respectively). Their experimental characterization shows an AFM ground state [65, 39, 67] which has led authors to expect a similar behavior also for V8. A comparison of these systems with the ones studied here is left for future work.
III Theory and computation
In this section we present the theoretical framework and the technical details of our calculations. We first discuss the spin Hamiltonian considered in this work and then outline the methodology used to compute the exchange interactions between the magnetic centers along the rings from ab initio calculations.
III.1 Spin Hamiltonian
The spin Hamiltonian $\hat{H}_{ex}$ used in this work is of the Heisenberg type, with magnetic exchange couplings between first- and second-nearest-neighbor spins, while interactions between more distant neighbors turn out to be negligible. $\hat{H}_{ex}$ accounts for the coupling between localized moments through their scalar product:
$\hat{H}_{ex}=J^{(1)}\sum_{i}\hat{\mathbf{S}}_{i}\cdot\hat{\mathbf{S}}_{i+1}+J^{(2)}\sum_{i}\hat{\mathbf{S}}_{i}\cdot\hat{\mathbf{S}}_{i+2}.$ (1)
Fig. 3 offers a visual representation of the magnetic couplings between the TM centers along the molecule.
Figure 3: Exchange coupling parameters $J^{\left(1\right)}$ and $J^{\left(2\right)}$ between magnetic ions.
III.2 Exchange interactions
The magnetic couplings present in the spin Hamiltonian (Eq. 1) will be obtained from ab initio calculations based on DFT, [63] in the Kohn-Sham formulation [64]. While the full characterization of MNMs magnetic properties would require wavefunction-based quantum chemistry approaches, these methods are largely impractical here, due to the large size of our systems. DFT represents an optimal compromise between accuracy and computational costs, allowing direct access to the systems ground state energy and selected perturbations around it.
The magnetic couplings are evaluated from ab initio calculations, by fitting the DFT total energy of suitably chosen spin configurations with the predictions of the spin Hamiltonian. Besides widely used local and semilocal approximations to the exchange-correlation (xc) functional (such as LDA[63, 64] and GGA [68]) we use the popular DFT+U [69, 70] and the extended DFT+U+V [71, 72] approaches, necessary to improve the $d$ states localization and achieve a more accurate representation of the magnetic moments and their interactions.
Spin-resolved DFT calculations are performed within the collinear framework, where each electron is in an eigenstate of the $\sigma_{z}$ Pauli matrix and the electronic magnetization is a scalar function. In this simplified framework, all magnetic moments align parallel to the same spin quantization axis, independent of the crystalline axes orientation (a non-collinear generalization will be discussed in a future work).
Over the last decades several methods have been introduced to compute the effective exchange couplings from first principles, including those based on Green’s functions and linear-response theory [73, 74, 75], and the four-state method [76, 77, 78].
Here we follow a somewhat simpler approach based on the broken-symmetry DFT method (BS-DFT)  [79, 80, 81, 82, 83, 84, 85, 86, 87]. More specifically, we adopt the projected BS-DFT, a method which relies on the same assumptions of the four-states method [76, 77, 78]. The approach consists in mapping the DFT total energies of various spin configurations onto the expectation values of the Hamiltonian (Eq. 1) on states that resemble the ones achieved from DFT (i. e. single-determinants constructed from the appropriate eigenstates of $\hat{S}^{z}$ on each magnetic ion). In practice, if we indicate these states as |$\ket{\sigma_{1}^{\left(k\right)},\dots,\sigma_{8}^{\left(k\right)}}$ (where $\sigma_{i}^{\left(k\right)}$ is the eigenvalue of $\hat{S}^{z}$ on the $i^{th}$ magnetic site for the $k^{th}$ spin configuration) our approach corresponds to identifying $E_{DFT}^{(k)}$ with $\bra{\sigma_{1}^{\left(k\right)},\dots,\sigma_{8}^{\left(k\right)}}\hat{H}_{ex}\ket{\sigma_{1}^{\left(k\right)},\dots,\sigma_{8}^{\left(k\right)}}$. When a sufficiently large number of states is considered (typical pools of states include FM, AFM and some intermediate configurations), the effective exchange parameters of $\hat{H}_{ex}$ can be extracted from the best fit of the DFT total energies.
Figure 4: Non-equivalent magnetic configurations used to extract the exchange coupling parameters. a) and b) are the FM and AFM configurations, respectively.
Fig. 4 shows some of the spin configurations typically used in the calculation of the exchange parameters.
After having computed the magnetic couplings, it is worth stressing that, while the structure of the molecule possesses a D4 symmetry (due to the alternate modulation of ligands groups), the magnetic centers are found to be approximately equivalent and deviations from a D8 symmetry can be safely neglected both for the modulus of the localized spins and for the interactions between them. In addition, due to the annular geometry of the systems considered here, periodic boundary conditions are assumed on the site indexes: $i+8=i$.
III.3 Computational details
Density Functional Theory (DFT) calculations were performed using the publicly available, plane-wave pseudopotential package Quantum ESPRESSO (QE) [88, 89, 90]. Spin-resolved, Hubbard-corrected calculations were based on a collinear-spin formulation of DFT. Specifically, we compare results obtained from two most popular approximations for the exchange-correlation (xc) functional: the local density approximation (LDA) [63, 64] and the PBE flavor of the generalized-gradient approximation (GGA) [68]. The pseudopotentials of the TMs were obtained from the QE PSlibrary archive [91] [92] and were chosen to be Ultrasoft (US) and projected-augmented wave (PAW) for GGA and LDA, respectively. Kohn-Sham wavefunctions and charge density were expanded in plane-waves up to a kinetic-energy cutoffs of 80Ry and 640Ry, respectively, in all cases.
Due to the complexity of the real molecular systems, a simplification of the structures was necessary to make the calculations affordable. This was achieved using the widely adopted “hydrogen saturation” procedure [53, 41, 43], which consists in replacing terminal methyl (-CH3) groups with single H atoms. The procedure was sequentially applied twice, and the resulting structure can be observed in Fig. 5 (to be compared with Fig. 2).
Figure 5: Reduced molecular structure (see text for explanation).
Overall, the pivalic groups were replaced by -O2CH groups. As a result, our simulated systems have chemical formulas Cr8F8(O2CH)16 and V8F8(O2CH)16. Since the rings magnetic properties stem from the localized TM ions $d$-electrons, the reduction is expected to have negligible effects on the systems’ magnetic behavior. Both simplified structures were optimized using both GGA and LDA and differences between the equilibrium structures of Cr8 and V8 were found marginal. Calculations on isolated molecules were performed by placing them in empty simulation boxes with a distance between periodic replicas of $\approx$ 10 Å. Correspondingly we used a $\Gamma$-point sampling of the Brillouin zone.
In order to improve the $d$ states localization and achieve a better representation of magnetic properties, we also used the DFT+U and DFT+U+V schemes. One-site $U$ and inter-site $V$ Hubbard parameters were computed using linear-response theory [93] within the automatic approach introduced in Refs. [94, 95, 96, 97]. To speed-up this calculation, a linear model of the rings’ local structure was used, i.e. a chain containing only two magnetic centers (instead of eight) with the same octahedral environment of the TM centers as in the whole molecule.
Figure 6: Linearized chain model of the ring molecule, used to compute the Hubbard parameters. The vertical lines delimitate the two-magnetic-sites unit cell (see text).
The chain is obtained from the simplified ring in Fig. 5 by changing the position of the ligand groups around the magnetic centers in such a way they can be arranged in a linear conformation. In the process, the structure of the local O-F octahedra around the TMs is maintained as close as possible to the original one. Test calculations of the Hubbard parameters in the ring (optimized with GGA) and in the chain yielded almost coincident results  [^1] A more exhaustive discussion about this procedure (and the importance of a comparison between curved and linear molecules) will be given in a separate publication.
In all cases the calculation of these parameters was performed self-consistently with respect to the electronic state (i.e. $U$’s and $V$’s were obtained perturbing a DFT+U(+V) ground state until input and output values converged). The obtained values are reported in Table 1, showing, for both systems, the considered xc functionals and flavours of the Hubbard correction (DFT+U and DFT+U+V), the value of the computed Hubbard parameters.
Cr8 V8
DFT+$U$ DFT+$U+V$ DFT+$U$ DFT+$U+V$
LDA $U$ 5.26 Cr-Cr 6.28 4.62 V-V 5.66
$V$ Cr-F 1.54 V-F 1.48
Cr-O 1.40 V-O 1.37
GGA $U$ 5.13 Cr-Cr 6.11 4.51 V-V 5.53
$V$ Cr-F 1.48 V-F 1.45
Cr-O 1.37 V-O 1.37
Table 1: Calculated Hubbard parameters (in eV).
It is important to remark that the results shown in Table 1 refer to the ground-state spin configuration of each system. However, test calculations have shown that the Hubbard parameters’ dependence on the arrangement of magnetic moments can be safely neglected. Table 1 shows relatively minor differences (0.1 - 0.2 eV) in the value of $U$ and $V$ between LDA and GGA, independently of the flavor of the Hubbard correction. In contrast, turning on the inter-site interaction ($V$) leads to a significant upward shift ($\approx$ 1 eV) of the on-site Hubbard $U$. At the same time the intersite $V$ exhibits surprisingly similar values for Cr and V, probably due to the close resemblance of the ligands’ structure around the TMs in the two systems. The on-site Hubbard parameters show instead a more pronounced difference between the two systems with those of Cr $\approx$ 0.6 eV higher than those of V. Finally, LDA yields higher values than GGA; the shift (non uniform and slightly more pronounced for $U$ than for $V$) is of the order of 0.1 eV.
IV Results
The main objective of this work is to evaluate the superexchange couplings between the magnetic sites of the considered molecules. Since we focus on collinear-spin calculations, only the isotropic part of the effective interactions are obtained  [^2] Higher-order terms will be evaluated and discussed in a future publication, also employing a non-collinear-spin DFT. Within these settings the expectation values of $\hat{H}_{ex}$ on the collinear spin states corresponding to the ones computed with DFT reduce to the energy of a classical Ising model:
$E^{\left(k\right)}=$ $\,J^{\left(1\right)}\sum_{i}\sigma_{i}^{\left(k\right)}\sigma_{i+1}^{\left(k\right)}+\,J^{\left(2\right)}\sum_{i}\sigma_{i}^{\left(k\right)}\sigma_{i+2}^{\left(k\right)}+c,$ (2)
where the additive constant $c$ fixes the energy reference. The considered spin configurations are identified by the sign of the spin on each magnetic site (the modulus being always maximum: $|\sigma_{i}|=S_{i}$) which is controlled by the proper initialization of the local magnetization. Typically, between the FM and the AFM states we consider five/six intermediate configurations (some of which are exemplified in Fig. 4) featuring various numbers of spin-flips in various positions.
A first rough estimate of the nearest-neighbor couplings can be also obtained from the direct comparison of the FM and AFM spin arrangements (see Fig. 4). The results for LDA and GGA are reported in Table 2, showing
Cr8 V8
${\rm J}_{LDA}$ 6.326 -38.370
${\rm J}_{GGA}$ 3.498 -3.434
Table 2: Nearest-neighbor exchange parameters (in meV) extracted from FM and AFM states.
that both functionals yield quite large exchange couplings [^3] We stress that the results shown in Table 2 are rather sensitive to the molecule’s structure (i.e., can change significantly for minor readjustments of the TM centers’ coordination shell) which, in turn, is often hard to optimize. This sensitivity disappears when using DFT+U or DFT+U+V. We thus believe it to be another effect of the overdelocalization of valence $d$ electrons in LDA or GGA.. In the case of Cr8, the values strongly overestimate the experimental one ($\sim$1.46 meV, from inelastic neutron scattering (INS) data [58] and electron paramagnetic resonance (EPR) measurements [38, 36]). The mismatch is larger in LDA, which for V8 yields an unrealistic $\sim$-38 meV value. Besides these quantitative inaccuracies, it is important to notice the difference in the sign of the couplings between V8 and Cr8. In agreement with literature [57, 38, 36, 58, 43, 52, 59, 53, 54] Cr8 is found to be AFM, driven by positive nearest-neighbor exchange interactions $J^{(1)}$. In contrast, a completely different picture emerges for V8, exhibiting a negative $J^{\left(1\right)}$ that stabilizes a FM order.
Importantly, using Hubbard corrections strongly modifies the exchange interactions. In the following we present the values obtained with DFT+U and DFT+U+V (DFT is either LDA or GGA) for both systems. In this case the fit on DFT total energy was based on a larger number of spin configurations (see Fig. 4 for a selection) which also gave access to the exchange couplings between second nearest neighbors (those between further neighbors were found negligible in all cases).
Functional $J$ Cr8 V8
LDA+U $J^{\left(1\right)}$ 0.843 -0.917
$J^{\left(2\right)}$0.013 0.231
GGA+U $J^{\left(1\right)}$ 0.628 -0.937
$J^{\left(2\right)}$0.011 0.098
Table 3: Cr8 and V8 exchange couplings (in meV) from collinear LDA+U and GGA+U calculations.
Table 3 collects the couplings obtained from LDA+U and GGA+U (on-site corrections only). While the couplings upon the Hubbard correction are hugely renormalized as compared to LDA and GGA, the signs of the nearest neighbor couplings are not affected, meaning that the character of the ground state in both systems remains the same: AFM for Cr8 ($J^{\left(1\right)}$ $>$ 0) and FM for V8 ($J^{\left(1\right)}$ $<$ 0). The values of $J^{\left(1\right)}$ are now of the expected order of magnitude and are much closer to the experimental value for Cr8 ($\sim$1.46 meV [58, 36, 38]). Both LDA+U and GGA+U show somewhat underestimated values.
Moreover, besides a FM $J^{\left(1\right)}$, V8 also features a sizeable second-nearest neighbor coupling $J^{\left(2\right)}$, that is positive and, in magnitude, represents a significant fraction of $J^{\left(1\right)}$. In fact, a sizeable $J^{\left(2\right)}$ is needed in order to fit the DFT total energies and make the system ”Heisenberg-like” (i.e., with couplings independent from the spin configuration). $J^{\left(2\right)}$ is also present for Cr8 but is one order of magnitude smaller than for V8. The presence of a relatively large and positive $J^{\left(2\right)}$ is an interesting feature of V8: competing (opposite sign) magnetic interactions of comparable strength might induce frustration and/or enhance the system’s susceptibility to external fields, enabling easier switching between different magnetic orders, an aspect that could be exploited in some applications.
In order to fully appreciate the role of the Hubbard corrections we also evaluated the couplings within the generalized DFT+U+V scheme [71]. The inter-site effective interaction ($V$) between electrons localized on neighboring sites is likely to influence the couplings between localized moments, typically mediated by anions’ extended orbitals between them. The results are shown in Table 4.
Functional $J$ Cr8 V8
LDA+U+V $J^{\left(1\right)}$ 1.199 -0.577
$J^{\left(2\right)}$0.014 0.248
GGA+U+V $J^{\left(1\right)}$ 0.908 -0.762
$J^{\left(2\right)}$0.011 0.174
Table 4: Cr8 and V8 exchange couplings (in meV) from collinear LDA+U+V and GGA+U+V calculations.
The comparison with Table 3 results clarifies immediately that the coupling $V$ (between the TMs and the anions in their first shell of neighbors) generally shifts the value of couplings upward (negative ones become less negative). This slightly improves the agreement for Cr8 with experiments compared to the DFT+U case. The shift is quantitatively more consistent for $J^{\left(1\right)}$ than $J^{\left(2\right)}$ and slightly more pronounced for LDA than GGA. Aside from these differences, however, the overall picture remains unchanged: positive (AFM) $J^{\left(1\right)}$’s and negligible $J^{\left(2\right)}$ for Cr8; negative (FM) $J^{\left(1\right)}$ and consistent, positive $J^{\left(2\right)}$ for V8.
In spite of their structural and chemical similarity (Cr and V are right next to each-other on the periodic table) the two molecules exhibit radically different ground states with an AFM order for Cr8 (mainly stabilized by $J^{\left(1\right)}$) and a FM arrangement for V8 that results from the competition between $J^{\left(1\right)}$ (FM, and numerically prevalent) and $J^{\left(2\right)}$ (AFM).
V Discussion
The results from the previous Section carry several messages that we can now examine in some detail.
First, as well known, LDA and GGA functionals are inadequate for accurately capturing the d-states localization. The consequences of this inaccuracy are dramatic for the properties studied in this work. In fact, given the typical small values of the interatomic exchange parameters, even a slight electron delocalization might lead to a gross overestimate of the couplings, as shown in Table 2. This traces back to the physics of superexchange, which mediates magnetic interactions between localized moments via intermediate anions with fully occupied orbitals. In fact, in a simple Hubbard model of the electrons responsible for the localized moments, the magnetic coupling, to the lowest order of perturbation theory, is proportional to $t^{2}/U$. Here $t$ represents the hopping parameter (or bandwidth), which quantifies the energetic advantage of forming bonds or crystalline itinerant states, while $U$ is the short-range on-site Coulomb repulsion. Clearly, overestimating $t$ and underestimating $U$, both consequences of the local/semilocal functionals, leads to severely off-scale magnetic interactions. This interpretation also explains the significant improvement in evaluating the exchange couplings achieved with DFT+U, which incorporates corrections to local/semilocal functionals inspired by the Hubbard model. By using LDA+U and GGA+U, we were able to predict $J^{\left(1\right)}$ for Cr8 with much closer agreement to the experimental value compared to LDA and GGA and, in general, obtain exchange couplings of the correct order of magnitude.
The agreement with experimental results on Cr8’s $J^{\left(1\right)}$ is further enhanced by using the extended DFT+U+V scheme with on-site and inter-site interactions (between TM magnetic centers and neighbor anions). While this result probably confirms the role of the intermediate anions orbitals in the superexchange mechanism, the overall qualitative picture is the same as with DFT+U.
Regarding the comparison between LDA and GGA, based on the value of $J^{\left(1\right)}$ in Cr8, LDA produces slightly more accurate results than GGA when corrected with some flavor of the Hubbard functional. While this result does not include the effects of structural relaxation, we can say that, again, the overall qualitative picture emerging from the two families of functionals is the same. However, in the perspective of performing non-collinear spin calculations, GGA xc kernels are sometimes reported to be more problematic than LDA [101]. Thus, LDA might be slightly preferable, especially when using a basic non-collinear-spin implementation of DFT.
From a physical point of view, the most important outcome to remark is the different behavior of Cr8 and V8. In spite of the identical ligand struture and of their chemical similarity, independently also from the specific functional used to model their behavior, the molecules exhibit a radically different ground-state magnetic order: AFM for Cr8, FM for V8. In addition, while the energetics of Cr8 can be adequately described using only the nearest-neighbor exchange interaction $J^{\left(1\right)}$, V8 also exhibits a significant next-nearest-neighbor coupling, $J^{\left(2\right)}$, with a magnitude comparable to $J^{\left(1\right)}$ but of opposite sign.
This behavior of V8 is rather unusual. Our preliminary calculations, where V centers are substituted with other TMs, consistently stabilize an AFM ground state. The V8 peculiarity seems to question the validity of the superexchange theory (also challenged by the presence of multiple anionic links between neighbor TM centers) and of the Goodenough-Kanamori-Anderson rules. However, it is important to note that in their simpler form, superexchange mechanisms are often referred to half-filling conditions. These mechanisms are derived from second-order perturbation theory applied to an Hubbard model with one orbital and one electron per site (half-filling). This condition is not satisfied by V8 that, in its 3+ oxidation state, hosts the two outermost electrons in a group of three quasi-degenerate $d$ states (this would be the $t_{2g}$ multiplet if the ligands’ octahedron around it had a perfect cubic symmetry). We believe that this less-than-half-filling condition of V3+ is what makes V8 to be FM. It is quite well established in literature that at less-than-half filling, a strong enough on-site repulsion between electrons ($U$ or $U-J$ for parallel spins) or, equivalently, a weak hopping amplitude ($t$), can cause a FM coupling between neighboring spins, mediated by the intermediate anion [102, 103].
To directly verify this point, we devised a simple Hubbard model with less-than-half-filled orbitals on the metallic centers, which we solved exactly to study the ground state energy as a function of $J/t$ ($J$ being the Hund’s exchange coupling for electrons on the same atomic site).
Figure 7: A schematic representation of the simple model adopted to study the superexchange coupling between less-than-half occupied metallic sites (M), as mediated by anions between them. $t$ is the hopping amplitude between metallic and anionic states; $\epsilon$ is the site energy of anionic states; $U$ and $U-J$ represent the on-site repulsion between electrons on the same metal with antiparallel or parallel spin, respectively.
As sketched in Fig. 7, the model contains two lateral transition-metal sites (indicated with M, each contributing two states) and two anionic sites (A) between them, each contributing one state. The total number of electrons is six, with anions levels guaranteed to be doubly-occupied by a negative site energy $\epsilon$, so that only one electron is present on each metallic site. This realizes a quarter-filling condition. The repulsion between electrons on the same metallic site is the Hubbard $U$possibly corrected by the exchange $J$ ($U$ or $U-J$ for antiparallel and parallel spins, respectively). The hybridization between metallic and anionic states is promoted by the hopping parameter $t$. The resulting 6-states Hubbard Hamiltonian can then be diagonalized exactly, e.g., using the HPHI code [104, 105].
Figure 8: Energy difference between the triplet and singlet ground states of the less-than-half-filled Hubbard model described in the text. All quantities are in units of the hopping amplitude $t$.
[…]
This observation confirms the idea that, for strong enough on-site exchange interactions, less-than-half-filled metallic states can couple ferromagnetically, even for geometries that would otherwise promote AFM configurations. Therefore, the anomalous ground state of V8 can, at least in part, be traced back to the lower filling of its $d$ states.
VI Conclusions
In this work we have presented a computational study of two ring-shaped molecular nano-magnets with eight magnetic sites, named Cr8 and V8. Fitting the DFT energies corresponding to various collinear-spin configurations we were able to extract the inter-site exchange magnetic couplings between the localized spins, using a number of different exchange-correlation functionals and Hubbard-based corrective schemes. Consistent with previous studies, we find that standard LDA/GGA DFT approximations fail to localize valence electrons on the $d$ states of the transition-metal magnetic centers, resulting in gross overestimations of the magnetic couplings. Energy corrections introduced with DFT+U and DFT+U+V schemes lead to significant improvements, with the inter-site coupling $V$ further refining the agreement with the experimental measurements. Independently of the functional used, and in agreement with previous investigations, Cr8 is found to have an antiferromagnetic ground state, with magnetic interactions limited to nearest-neighbors. On the other hand, V8 is found to have a ferromagnetic ground state with nearest-neighbor interactions competing with antiferromagnetic next-nearest-neighbor interactions of the same order of magnitude.
The anomalous behavior of V8 can be rationalized based on the less-than-half filling condition of the $d$ states submanifold hosting the valence electrons responsible for the localized moments, as qualitatively demonstrated by the exact diagonalization of a simple Hubbard model mimicking this situation.
VII Acknowledgements
B. M. M. and M. C. acknowledge support from NQSTI within the PNRR MUR project PE0000023-NQSTI. Part of the calculations of this work were realized thanks to the high-performance computing resources and support made available by CINECA (through awards within the ISCRA initiative). Moreover, via A.F. membership of the UK’s HEC Materials Chemistry Consortium, funded by EPSRC (EP/X035859), this work also used the ARCHER2 UK National Supercomputing Service (http://www.archer2.ac.uk).